OpenStudy (anonymous):

If a and b are two sides of a right triangle, find the length of the hypotenuse, c. a=6, b=2, c=?

OpenStudy (mertsj):

\[a^2+b^2=c^2\]Try the Pythagorean Theorem:

OpenStudy (saifoo.khan):

Where "c" is the hypotenuse.

OpenStudy (saifoo.khan):

\[6^2 + 2^2 = c^2\] \[36 + 4 = c^2 \] \[c = \sqrt{40}\] \[c = 2 \sqrt{10}\]